you can mow 10 square yards of grass in one minute. how long does it take you to mow a triangular plot with height 25 yards and base 24 yards? how long does it take you to mow a rectangular plot with base 24 yards and height 36 yards?

Answer by checkley77(12844)

1/10=x/[(25*24)/2]
1/10=x/(25*12)
1/10=X/300
10X=300
X=300/10
X=30 MINUTES TO MOW THE TRIANGLE.
----------------------------
1/10=X/(24*36)
1/10=X/864
10X=864
X=864/10
X=86.4 MINUTES TO MOW THE RECTANGLE.